# Dynamical signatures of the Yang-Lee edge singularity in non-Hermitian systems

### v1

@@ I Introduction
Yang-Lee edge singularities are fascinating nonunitary critical points with simple unconventional universality classes that depend solely on the dimensionality and can be described by a nonunitary CFT [5, 6, 7]. However, the experimental realization of Yang-Lee edge singularities is particularly challenging as they appear in complex physical parameters. One proposal to realize the effect of complex physical parameters is to use a probe spin coupled to a system, where the Lee-Yang zeros correspond to the zeros of spin coherences [8]. Recently, the Lee-Yang zeros have been observed by directly measuring the coherence of a probe spin [9], in contrast to previous indirect detection experiment [10, 11]. Yang-Lee edge singularities, regarded as exceptional points [12, 13, 14, 15, 16, 17], represent nonunitary criticality in non-Hermitian systems. Thanks to the development of non-Hermitian physics [18, 19, 20, 21, 22, 23, 24, 25, 26, 27], Yang-Lee edge singularities have recently been observed in non-Hermitian quantum systems [28, 29]. This observation, in accordance with quantum-classical correspondence [30], has led to the derivation of critical exponents for the one-dimensional classical Ising model. Despite proposals for realizing Yang-Lee edge singularities in one-dimensional quantum many-body systems, such as utilizing Kibble-Zurek scaling [31], exploiting the quantum coherence of a probe spin [32], and investigating dissipative Rydberg atomic arrays [33], achieving the realization of the critical points and critical exponents of Yang-Lee edge singularities remains elusive.
@@
In this paper, we will use the LEs, which are readily achievable in non-Hermitian experiments [28, 46], to detect the critical points and critical exponents of Yang-Lee edge singularities, which are understood as exceptional points. We implement biorthogonal LEs in the associated bases and study their responses in both the non-Hermitian transverse field Ising model and the Yang-Lee model. Our analysis reveals that the Yang-Lee edge singularity can be characterized by the dynamical signatures of different phases, offering a direct observation of Yang-Lee edge singularities in non-Hermitian many-body systems.
This paper is organized as follows. In Sec.II, we introduce both the concept of the self-normal Loschmidt echo and the associated-biorthogonal Loschmidt echo. In Sec.III, we present the results for the non-Hermitian transversed field Ising model. In Sec.IV, we introduce and analyze phase transitions in the Yang-Lee model. In Sec.V, we summarize our results.
Figure 1: Schematic diagram of Loschmidt echo. (a) The self-normal Loschmidt echo $L_{S}$, (b) The associated biorthogonal Loschmidt echo $L_{A}$. Blue symbols represents the right eigenstates, and yellow symbols denotes their associated states.
@@ II Loschmidt echoes
where $|\psi_{n}(\lambda_{i})\rangle$ denotes the initial state, which is the $n$th eigenstate of the Hamiltonian $H(\lambda)$ at parameter $\lambda_{i}$, and $|\psi_{n}(\lambda_{f},\lambda_{i}),t\rangle$ represents the time-evolved state of $|\psi_{n}(\lambda_{i})\rangle$ following a quench from $H(\lambda_{i})$ to $H(\lambda_{f})$.
@@
where $\epsilon_{n}(\lambda)$ and $\epsilon^{\ast}_{n}(\lambda)$ represent right and left eigenvalues. It is known that biorthogonal eigenvectors satisfy both the completeness relation [44],
@@
The time-evolved states in biorthogonal bases under a quench from $\lambda_{i}$ to $\lambda_{f}$ are given by [39],
$|\psi^{R}(\lambda_{f},\lambda_{i},t)\rangle=e^{-iH_{f}t}|\psi^{R}(\lambda_{i})\rangle,$ (6)
$|\psi^{L}(\lambda_{f},\lambda_{i},t)\rangle=e^{-iH_{f}^{\dagger}t}|\psi^{L}(\lambda_{i})\rangle.$ (7)
@@
$H_{f}(\lambda)=H_{i}(\lambda)+\delta H(\lambda).$ (8)
@@
$L_{S}(\lambda_{f},\lambda_{i},t)=\langle\psi^{R}(\lambda_{i})|\psi^{R}(\lambda_{f},\lambda_{i},t)\rangle\langle\psi^{R}(\lambda_{f},\lambda_{i},t)|\psi^{R}(\lambda_{i})\rangle,$ (9)
using only right eigenstates $|\psi^{R}_{n}(\lambda)\rangle$, and the biorthogonal LE,
$L_{B}(\lambda_{f},\lambda_{i},t)=\langle\psi^{L}(\lambda_{i})|\psi^{R}(\lambda_{f},\lambda_{i},t)\rangle\langle\psi^{L}(\lambda_{f},\lambda_{i},t)|\psi^{R}(\lambda_{i})\rangle.$ (10)
using both right eigenstates $|\psi^{R}_{n}(\lambda)\rangle$ and left eigenstates $|\psi^{L}_{n}(\lambda)\rangle$.
It is expected that the biorthogonal LE should describe the phase transition [39], as a non-Hermitian Hamiltonian is diagonalized in biorthogonal bases. Indeed, it does describe phase transitions in the real-energy regime. However, in the complex-energy regime, the biorthogonal LE, $L_{B}(\lambda_{f},\lambda_{i},t)$, becomes meaningless. This is because the definition of the biorthogonal LE in Eq.(10) is unitary, whereas a non-Hermitian system evolves non-unitarily in the complex-energy regime.
In order to describe proper evolutions of the biorthogonal LE, we introduce associated states as outlined in Ref.[44, 45]. For any arbitrary right eigenstate $|\Psi^{R}\rangle$, we consider it as a superposition of a series of eigenstates and define the associated state $|\tilde{\Psi}^{L}\rangle$ using the operation,
$|\Psi^{R}\rangle=\sum_{n}c_{n}|\psi_{n}^{R}\rangle\rightarrow|\tilde{\Psi}^{L}\rangle=\sum_{n}c_{n}|\psi_{n}^{L}\rangle,$ (11)
where the coefficient $c_{n}=\langle\psi^{L}_{n}|\Psi^{R}\rangle$ is the same as in $\langle\tilde{\Psi}^{L}|=\sum_{n}c_{n}^{\ast}\langle\psi^{L}_{n}|$. Here, $|\tilde{\Psi}^{L}\rangle$ denotes the associated state.
As a result, the biorthogonal LE in the associated state bases, dubbed as the associated-biorthogonal LE, is defined as [c.f. Figs.1(b)]:
$L_{A}(\lambda_{f},\lambda_{i},t)=\langle\tilde{\psi}^{L}(\lambda_{i})|\psi^{R}(\lambda_{f},\lambda_{i},t)\rangle\langle\tilde{\psi}^{L}(\lambda_{f},\lambda_{i},t)|\psi^{R}(\lambda_{i})\rangle.$ (12)
@@ III Non-Hermitian transversed field Ising model
where a complex field is introduced along the $y$-direction. In this non-Hermitian transversed field Ising (NHTI) model, $J\geq 0$ denotes the strength of interaction between adjacent spins in the $x$-direction, defining the ferromagnetic coupling. $\lambda$ denotes the strength of the transverse field, whereas $\gamma$ introduces an additional field that quantifies the non-Hermiticity in the system. Here, $\sigma_{j}^{x},\sigma_{j}^{y},\sigma_{j}^{z}$ denote there Pauli matrices along $x,y,z$-directions at lattice site $j$. Throughout the paper, we impose periodic boundary conditions such that $\sigma_{L+1}^{x,y,z}=\sigma_{1}^{x,y,z}$, where $N$ represents the length of the chain.
Figure 3: Time evolutions of LEs of the Yang-Lee model. (a)(b) Self-normal and associated-biorthogonal LEs for $N=12$, $\lambda=0.6$ with respect to $h$ and $t$. Here, $\delta t=1$ and $\delta h=0.001$ are used. (c)(d) Self-normal and associated-biorthogonal LEs at $t=40$, where the LEs exhibit a change around $h=0.104$, signaling a phase transition. Through the critical point, $h=0.104$, the system undergoes a transition from the $\mathcal{PT}$ symmetric phase to the $\mathcal{PT}$-symmetry broken phase.
The NHTI model can be solved exactly using a similarity transformation [39, 47, 48, 49, 50]. It is known that this model exhibits an exceptional point at $\gamma=\pm 1$ [39, 47, 48, 49, 50]. To determine if the LEs can capture this phase transition, we compute both the self-normalized LE and the associated biorthogonal LE in the NHTI model. We use the equation (2) to obtain right eigenstates varying $\gamma$ from $-2$ to $2$ at $N=8$, $J=1$ and $\lambda=0.5$ during the simulations. We identify the right eigenstate with the lowest real part of the energy as the ground state, $|\psi^{R}_{0}\rangle$. The time-evolved state, $|\psi^{R}_{0}(t)\rangle$ is then obtained using Equation (6). The associated states $|\tilde{\psi}^{L}_{0}\rangle,|\tilde{\psi}^{L}_{0}(t)\rangle$, used for calculating the associated-biorthogonal LE, are derived from Equation (11).
@@
Table 1: Critical points for finite systems with $\lambda$ ranging from $0.1$ to $0.9$. Here $h_{N}$ represents the finite critical points obtained from the LEs, while $h_{\text{LE}}$ denotes the critical points we derived in the thermodynamic limit. The term MPRG refers to critical values derived using the MPRG method [16], and $h_{c}$ represents critical values obtained from exact diagonalization [13] and Loop-TNR methods [17].
$\lambda$ $h_{\text{N=6}}$ $h_{\text{N=8}}$ $h_{\text{N=10}}$ $h_{\text{N=12}}$ $h_{\text{LE}}$ MPRG $h_{c}$
0.1 0.642 0.639 0.638 0.637 0.6360 0.636 0.6366
0.2 0.469 0.464 0.461 0.460 0.4579 0.457 0.4585
0.3 0.345 0.338 0.334 0.333 0.3300 0.328 0.3300
0.4 0.252 0.242 0.237 0.236 0.2317 0.230 0.2320
0.5 0.181 0.169 0.163 0.161 0.1561 0.154 0.1562
0.6 0.128 0.114 0.107 0.104 0.0985 0.095 0.0981
0.7 0.088 0.074 0.067 0.063 0.0579 0.052 0.0548
0.8 0.060 0.046 0.039 0.034 0.0292 0.021 0.0247
0.9 0.041 0.027 0.021 0.017 0.0119 - 0.0065
@@ IV Yang-Lee Model
where $\lambda$ denotes the strength of the nearest-neighbor interaction along $x$-direction, and $h$ represents an external complex magnetic field, which can be interpreted as the contribution from dissipation acting on the system. This model undergoes a phase transition [12, 13, 14, 15, 16, 17], with the critical point referred to as the Yang-Lee edge singularity. In contrast to the NHTI model, the Yang-Lee model is a non-integrable model that cannot be solved exactly. Instead, it is analyzed using the framework of non-unitary conformal field theory (CFT) [5, 6, 7]. Additionally, numerical simulations [12, 13, 14, 15, 16, 17] have confirmed the results predicted by non-unitary CFT. In the following, we will focus on the dynamical description of the Yang-Lee edge singularity through LEs, which are more easily realized experimentally.
@@
In the Yang-Lee model, the Yang-Lee edge singularity is an exceptional point [12, 13, 14, 15, 16, 17]. Specifically, in the $\mathcal{PT}$ symmetric regime, the eigenvalues are real. Conversely, in the $\mathcal{PT}$-symmetry broken regime, the eigenvalues become complex and appear in conjugate pairs. This should lead to distinct dynamical behaviors, as demonstrated in the NHTI model. To evaluate the validity of this argument, we compute both the self-normal LE and the associated-biorthogonal LE for $N=12$ at $\lambda=0.6$ with $h$ ranging from $0.08$ to $0.13$. The results of the LEs are presented in Figs.3. We observe that, similar to the NHTI model, the associated-biorthogonal LE decays rapidly from an initial value of one to zero, while the self-normal LE exhibits an approximately linear decay. Critical points are determined by the intersection of the linear decay line with $L_{S}=1$ and the point closest to $L_{A}=1/2$, respectively. Interestingly, both LE methods indicate the critical point $h_{N}=0.104$.
@@
Figure 4: (a) Phase diagram of the Yang-Lee model as functions of $\lambda$ and $h$. In this phase diagram, the red dot denotes the critical values $h_{\text{LE}}$ obtained from the LEs. The yellow region signifies the $\mathcal{PT}$-symmetric phase, while the purple region represents the $\mathcal{PT}$-symmetry broken phase. The critical line between these regions marks the Yang-Lee singularity edge. (b) Finite-size scaling of critical values $h_{N}$ with respect to the reciprocal of lattice size $N$ at $\lambda=0.6$. Here, we use $\beta=2.4$ for fitting.
@@
We note that the critical point $h_{N}=0.104$ for the $N=12$ lattice sites is slightly larger than the critical value $h_{c}=0.0981$ in the thermodynamic limit [13, 17]. To more accurately locate the Yang-Lee edge singularity, we perform simulations for both LEs with $N=4,6,8,10,12$ by varying $h$ and $\lambda$. The critical values obtained for each lattice size are presented in Table 1. The critical values in the thermodynamic limit are obtained from finite-size scaling, given by:
@@
where $h_{\text{LE}}$ is the critical value in the thermodynamic limit, and $\beta$ is the critical exponent, which is $\beta=12/5$ according to non-unitary CFT [5, 6, 7, 12, 13, 33]. The phase diagram of the Yang-Lee model is presented in Figs.4(a). Specifically, the fitting results for $\lambda=0.6$ are presented in Figs.4(b), which are consistent with the analytical result from non-unitary CFT and other numerical data [5, 6, 7, 12, 13, 14, 15, 16, 17, 33].
It is noteworthy that the critical values derived from the LEs for $\lambda<0.7$ more closely match the numerical critical points $h_{c}$ obtained from other methods, compared to the case where $\lambda>0.7$ (see Table 1). This discrepancy arises because, for $\lambda>0.7$, the critical point is nearer to the Ising transition, resulting in weaker non-Hermiticity. As a result, non-unitary time evolution is less sensitive in this regime compared to $\lambda<0.7$. Additionally, we note that it is observed that the associated-biorthogonal LE is more sensitive than the self-normal LE near the exceptional points.

### v2

@@ I Introduction
Yang-Lee edge singularities are fascinating nonunitary critical points with simple unconventional universality classes that depend solely on the dimensionality and can be described by a nonunitary CFT [5, 6, 7]. However, the experimental realization of Yang-Lee edge singularities is particularly challenging as they appear in complex physical parameters. One proposal to realize the effect of complex physical parameters is to use a probe spin coupled to a system, where the Lee-Yang zeros correspond to the zeros of spin coherences [8]. Recently, the Lee-Yang zeros have been observed by directly measuring the coherence of a probe spin [9], in contrast to previous indirect detection experiment [10, 11]. Yang-Lee edge singularities, regarded as exceptional points [12, 13, 14, 15, 16, 17], represent nonunitary criticality in non-Hermitian systems. Thanks to the development of non-Hermitian physics [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], Yang-Lee edge singularities have recently been observed in non-Hermitian quantum systems [29, 30]. This observation, in accordance with quantum-classical correspondence [31], has led to the derivation of critical exponents for the one-dimensional classical Ising model. Despite proposals for realizing Yang-Lee edge singularities in one-dimensional quantum many-body systems, such as utilizing Kibble-Zurek scaling [32], exploiting the quantum coherence of a probe spin [33], and investigating dissipative Rydberg atomic arrays [34], achieving the realization of the critical points and critical exponents of Yang-Lee edge singularities remains elusive.
@@
In this paper, we will use the LEs, which are readily achievable in experiments [47, 48, 49], to detect the critical phenomena [50] and critical exponents of Yang-Lee edge singularities [29] understood as exceptional points. We implement biorthogonal LEs in the associated bases and study their responses in both the non-Hermitian transverse field Ising model and the Yang-Lee model. Our analysis reveals that the Yang-Lee edge singularity can be characterized by the dynamical signatures of different phases, offering a direct observation of Yang-Lee edge singularities in non-Hermitian many-body systems.
This paper is organized as follows. In Sec.II, we introduce the concepts of the self-normal Loschmidt echo and the associated-biorthogonal Loschmidt echo. In Sec.III, we present the results for the non-Hermitian transversed field Ising model. In Sec.IV, we introduce and analyze phase transitions in the Yang-Lee model. In Sec.V, we summarize our results.
Figure 1: Schematic diagram of Loschmidt echoes. (a) The self-normal Loschmidt echo $L_{S}$, (b) The associated biorthogonal Loschmidt echo $L_{A}$. Blue symbols represents the right eigenstates, and yellow symbols denotes their associated states.
@@ II Loschmidt echoes
where $|\psi_{n}(\lambda_{i})\rangle$ denotes the initial state, which is the $n$th eigenstate of the Hamiltonian $H(\lambda)$ at parameter $\lambda_{i}$, and $|\psi_{n}(\lambda_{f},\lambda_{i},t)\rangle$ represents the time-evolved state of $|\psi_{n}(\lambda_{i})\rangle$ following a quench from $H(\lambda_{i})$ to $H(\lambda_{f})$.
@@
where $\epsilon_{n}(\lambda)$ and $\epsilon^{\ast}_{n}(\lambda)$ represent the $n$th right and left eigenvalues. It is known that biorthogonal eigenvectors satisfy both the completeness relation [45],
@@
The time-evolved states of $|\psi^{R}_{0}(\lambda_{i})\rangle$ and $|\psi^{L}_{0}(\lambda_{i})\rangle$ in biorthogonal bases under a quench from $\lambda_{i}$ to $\lambda_{f}$ are given by [40],
$|\psi^{R}_{0}(\lambda_{f},\lambda_{i},t)\rangle=e^{-iH_{f}t}|\psi^{R}_{0}(\lambda_{i})\rangle,$ (6)
$|\psi^{L}_{0}(\lambda_{f},\lambda_{i},t)\rangle=e^{-iH_{f}^{\dagger}t}|\psi^{L}_{0}(\lambda_{i})\rangle.$ (7)
@@
$H(\lambda_{f})=H(\lambda_{i})+\delta H(\lambda).$ (8)
@@
$L_{S}(\lambda_{f},\lambda_{i},t)=\langle\psi^{R}_{0}(\lambda_{i})|\psi^{R}_{0}(\lambda_{f},\lambda_{i},t)\rangle\langle\psi^{R}_{0}(\lambda_{f},\lambda_{i},t)|\psi^{R}_{0}(\lambda_{i})\rangle,$ (9)
using only right eigenstates $|\psi^{R}_{0}(\lambda)\rangle$, and the biorthogonal LE,
$L_{B}(\lambda_{f},\lambda_{i},t)=\langle\psi^{L}_{0}(\lambda_{i})|\psi^{R}_{0}(\lambda_{f},\lambda_{i},t)\rangle\langle\psi^{L}_{0}(\lambda_{f},\lambda_{i},t)|\psi^{R}_{0}(\lambda_{i})\rangle.$ (10)
using both right eigenstates $|\psi^{R}_{0}(\lambda)\rangle$ and left eigenstates $|\psi^{L}_{0}(\lambda)\rangle$.
It is expected that the biorthogonal LE should describe the phase transition [40], as a non-Hermitian Hamiltonian is diagonalized in biorthogonal bases. In fact, it does describe phase transitions in the real-energy regime [40]. However, in the complex-energy regime, the biorthogonal LE, $L_{B}(\lambda_{f},\lambda_{i},t)$, becomes meaningless. This is because the biorthogonal LE, as defined in Eq.(10), is unitary, whereas a non-Hermitian system evolves non-unitarily in the complex-energy regime.
In order to describe time evolution of the LE in the complex-energy regime within the biorthogonal bases, we introduce associated states as outlined in Ref.[45, 46]. For a right eigenstate $|\psi^{R}_{0}\rangle$, we consider it as a superposition of a series of eigenstates and define the associated state $|\tilde{\psi}^{L}_{0}\rangle$ using the operation,
$|\psi^{R}_{0}\rangle=\sum_{n}c_{n}|\phi_{n}^{R}\rangle\rightarrow|\tilde{\psi}^{L}_{0}\rangle=\sum_{n}c_{n}|\phi_{n}^{L}\rangle,$ (11)
where the coefficient $c_{n}=\langle\phi^{L}_{n}|\psi^{R}_{0}\rangle$ is the same as in $\langle\tilde{\psi}^{L}_{0}|=\sum_{n}c_{n}^{\ast}\langle\phi^{L}_{n}|$. Here, $|\phi_{n}^{R}\rangle$ and $|\phi_{n}^{L}\rangle$ forms a biorthogonal basis and $|\tilde{\psi}^{L}\rangle$ denotes the associated state.
The biorthogonal LE in the associated state bases, dubbed as the associated-biorthogonal LE, is defined as [c.f. Figs.1(b)]:
$L_{A}(\lambda_{f},\lambda_{i},t)=\langle\tilde{\psi}^{L}_{0}(\lambda_{i})|\psi^{R}_{0}(\lambda_{f},\lambda_{i},t)\rangle\langle\tilde{\psi}^{L}_{0}(\lambda_{f},\lambda_{i},t)|\psi^{R}_{0}(\lambda_{i})\rangle.$ (12)
@@ III Non-Hermitian transversed field Ising model
where a complex field is introduced along the $y$-direction. In this non-Hermitian transversed field Ising (NHTI) model, $J\geq 0$ denotes the strength of interaction between adjacent spins in the $x$-direction, defining the ferromagnetic coupling. $\lambda$ denotes the strength of the transverse field, whereas $\gamma$ introduces an additional field that quantifies the non-Hermiticity in the system. Here, $\sigma_{j}^{x},\sigma_{j}^{y},\sigma_{j}^{z}$ denote the Pauli matrices along $x,y,z$-directions at lattice site $j$. Throughout the paper, we impose periodic boundary conditions such that $\sigma_{L+1}^{x,y,z}=\sigma_{1}^{x,y,z}$, where $N$ represents the length of the chain.
Figure 3: Time evolutions of LEs of the Yang-Lee model. (a)(b) Self-normal and associated-biorthogonal LEs for $N=12$, $\lambda=0.6$ with respect to $h$ and $t$. Here, $\delta t=1$ and $\delta h=0.001$ are used. (c)(d) Self-normal and associated-biorthogonal LEs at $t=40$, where the LEs exhibit a change around $h_{N}=0.104$, signaling a phase transition. Here, the blue circle and red square symbols represent the numerical data without and with disorders (by averaging over $10$ disorder pertubations), respectively. The green dashed line in (c) denotes the linear fit, serving as a guide for the eye. Through the critical point, $h_{N}=0.104$, the system undergoes a transition from the $\mathcal{PT}$ symmetric phase to the $\mathcal{PT}$-symmetry broken phase.
The NHTI model can be solved exactly using a similarity transformation [40, 51, 52, 53, 54]. It is known that this model exhibits an exceptional point at $\gamma=\pm 1$ [40, 51, 52, 53, 54]. To determine if the LEs can capture this phase transition, we compute both the self-normal LE and the associated-biorthogonal LE in the NHTI model, where $\delta\lambda\equiv\delta\gamma$. We use equation (2) to obtain right eigenstates varying $\gamma$ from $-2$ to $2$ at $N=8$, $J=1$ and $\lambda=0.5$ and identify the right eigenstate with the lowest real part of the energy as the ground state, $|\psi^{R}_{0}\rangle$ during the simulations. The time-evolved state, $|\psi^{R}_{0}(t)\rangle$ is then obtained using Equation (6). The associated states $|\tilde{\psi}^{L}_{0}\rangle,|\tilde{\psi}^{L}_{0}(t)\rangle$, used for calculating the associated-biorthogonal LE, are derived from Equation (11).
@@
Table 1: Critical points of the Yang-Lee model for $\lambda$ ranging from $0.1$ to $0.9$ in the thermodynamic limit. Here $h_{\text{LE}}$, $h_{\text{MPRG}}$, $h_{\text{ED}}$ and $h_{\text{TNR}}$ represent the critical points $h_{c}$ obtained from the LEs, the matrix product renormalization group (MPRG) [16], the exact diagonalization (ED) [13] and the Loop tensor network renormalization (TNR) [17] methods, respectively.
[…]
0.1 0.6360 0.636 0.6366 0.6366
0.2 0.4579 0.457 0.4585 0.4585
0.3 0.3300 0.328 0.3300 0.3300
0.4 0.2317 0.230 0.2320 0.2320
0.5 0.1561 0.154 0.1562 0.1562
0.6 0.0985 0.095 0.0981 0.0981
0.7 0.0579 0.052 0.0548 0.0548
0.8 0.0292 0.021 0.0247 0.0247
0.9 0.0119 - 0.0065 0.0065
@@ IV Yang-Lee Model
where $\lambda$ denotes the strength of the nearest-neighbor interaction along $x$-direction, and $h$ represents an external complex magnetic field, which can be interpreted as the contribution from dissipation acting on the system. This model undergoes a phase transition [12, 13, 14, 15, 16, 17], with the critical point referred to as the Yang-Lee edge singularity. In contrast to the NHTI model, the Yang-Lee model is a non-integrable model that cannot be solved exactly. Instead, it is analyzed using the framework of non-unitary conformal field theory (CFT) [5, 6, 7]. Additionally, numerical simulations [12, 13, 14, 15, 16, 17] have confirmed the results predicted by non-unitary CFT. In the following, we will focus on the dynamical description of the Yang-Lee edge singularity through LEs, which are more easily realizable in experiments [47, 48, 49].
@@
In the Yang-Lee model, the Yang-Lee edge singularity is an exceptional point [12, 13, 14, 15, 16, 17]. Specifically, in the $\mathcal{PT}$ symmetric regime, the eigenvalues are real. Conversely, in the $\mathcal{PT}$-symmetry broken regime, the eigenvalues become complex and appear in conjugate pairs. This should lead to distinct dynamical behaviors, as demonstrated in the NHTI model. To evaluate the validity of this argument, we compute both the self-normal LE and the associated-biorthogonal LE with $h$ ranging from $0.08$ to $0.13$, using $\delta\lambda\equiv\delta h$ for $N=12$ at $\lambda=0.6$. The results of the LEs are presented in Figs.3. We observe that, similar to the NHTI model, the associated-biorthogonal LE decays rapidly from an initial value of one to zero, while the self-normal LE exhibits an approximately linear decay. Moreover, the behavior of the LEs is examined for robustness against random perturbations of the form $i\epsilon\sigma^{x}_{j}$, where $\epsilon$ is uniformly distributed within the interval $[-\delta h,\delta h]$ [c.f. Figs.3(c) and (d)]. Critical points are determined by the intersection of the linear decay line with $L_{S}=1$ and the point closest to $L_{A}=1/2$, respectively. Interestingly, both LE methods indicate the critical point $h_{N}=0.104$.
@@
Figure 4: (a) Phase diagram of the Yang-Lee model as functions of $\lambda$ and $h$. In this phase diagram, the red dot denotes the critical values $h_{\text{LE}}$ obtained from the LEs in the thermodynamic limit. The yellow region signifies the $\mathcal{PT}$-symmetric phase, while the purple region represents the $\mathcal{PT}$-symmetry broken phase. The solid line (served a guide for the eye) between these regions marks the Yang-Lee singularity edge. (b) Finite-size scaling of critical values $h_{N}$ with respect to the reciprocal of lattice size $N$ at $\lambda=0.6$. Here, the circles represent the critical values $h_{\text{N}}$ for finite lattice sizes, while the solid line is the fitting curve obtained using Eq.(15) with $\beta=12/5$, yielding the critical value shown in (a).
@@
However, we note that the critical point $h_{N}=0.104$ [see Appendix A] for the $N=12$ lattice sites is slightly larger than the critical value $h_{c}=0.0981$ in the thermodynamic limit [13, 17]. The difference arises from the finite-size effect. To more accurately locate the Yang-Lee edge singularity, we perform simulations for both LEs with $N=4,6,8,10,12$ by varying $h$ and $\lambda$. The critical values for each $N$ are presented in Table A1 in Appendix A. The critical values, $h_{\text{LE}}$, in the thermodynamic limit are obtained from finite-size scaling, given by:
@@
where $\beta$ is the critical exponent, which is $\beta=12/5$ according to non-unitary CFT [5, 6, 7, 12, 13, 34]. The phase diagram of the Yang-Lee model is presented in Figs.4(a). Specifically, the numerical data and the fitting curve for $\lambda=0.6$ are presented in Figs.4(b), which are consistent with the analytical result from non-unitary CFT and other numerical results [5, 6, 7, 12, 13, 14, 15, 16, 17, 34].
It is noteworthy that the critical values $h_{\text{LE}}$ derived from the LEs for $\lambda<0.7$ more closely match the numerical critical points $h_{c}$ obtained from other methods, compared to the case where $\lambda>0.7$ (see Table 1). This discrepancy arises because, for $\lambda>0.7$, the critical point is nearer to the Ising transition, resulting in weaker non-Hermiticity. As a result, non-unitary time evolution is less sensitive in this regime compared to $\lambda<0.7$. Additionally, it is observed that the associated-biorthogonal LE is more sensitive than the self-normal LE near the exceptional points. We note that it is feasible to compute the self-normal LEs using the time-evolving block decimation method [55] beyond unitary evolution [41]. However, numerically calculating the associated biorthogonal-biorthogonal LEs for large systems is more challenging, as determining the associated state requires the entire set of eigenvectors for constructing the projector.
Finally, we briefly discuss the experimental implementation of the self-normal LEs in the Yang-Lee model, following a protocol similar to the realization of dynamical quantum phase transitions in Hermitian systems [47]. First, the initial state is prepared as the ground state of the initial Hamiltonian $H_{0}$ at $h=h_{0}$. Second, the system is suddenly switched to the Hamiltonian $H_{f}$ with a small change $\delta h$, and it evolves according to Eq.(6). Third, the self-normal LEs, which represent the return probabilities to the ground state, are measured at discrete times. By varying the initial state with different values of $h_{0}$ and repeating these steps, experimental data can be obtained, as shown in Figs.3. The properties of the Yang-Lee edge singularity are then analyzed using finite-size scaling theory, as demonstrated in Figs.4(b).